Please help me put by giving me detailed step by step solution

Thank you

b) – In(x2 - 1)2 = 2In V2x2 – 3x+1
Answer #1

som n (22=132 = 2 m 0222-3341 Z ln (x²-1) = 21m (2x2-3x4132 In (x2-1) = ln (2x2-3x4157 a fn(x2-1) = Xm (2x²-3x+1) - 0²-1= 2x²